# Zero-Inflated Tweedie Boosted Trees with CatBoost for Insurance Loss Analytics

### v1

@@ Abstract
In this paper, we explore advanced modifications to the Tweedie regression model in order to address its limitations in modeling aggregate claims for various types of insurance such as automobile, health, and liability. Traditional Tweedie models, while effective in capturing the probability and magnitude of claims, usually fall short in accurately representing the large incidence of zero claims. Our recommended approach involves a refined modeling of the zero-claim process, together with the integration of boosting methods in order to help leverage an iterative process to enhance predictive accuracy. Despite the inherent slowdown in learning algorithms due to this iteration, several efficient implementation techniques that also help precise tuning of parameter like XGBoost, LightGBM, and CatBoost have emerged. Nonetheless, we chose to utilize CatBoost, a efficient boosting approach that effectively handles categorical and other special types of data. The core contribution of our paper is the assembly of separate modeling for zero claims and the application of tree-based boosting ensemble methods within a CatBoost framework, assuming that the inflated probability of zero is a function of the mean parameter. The efficacy of our enhanced Tweedie model is demonstrated through the application of an insurance telematics dataset, which presents the additional complexity of compositional feature variables. Our modeling results reveal a marked improvement in model performance, showcasing its potential to deliver more accurate predictions suitable for insurance claim analytics.
Keywords: Tweedie, zero-inflated models, tree-based boosting, gradient boosting, CatBoost, compositional features, aggregate claims models.
@@ 1 Introduction
Insurance loss analytics, which is the backbone of modern insurance risk management strategies, involves a systematic collection, processing, and analysis of observed insurance loss data to extract insights and produce predictive models to help quantify and manage risk exposure. Actuaries and risk managers have historically endured the use of data analytics to uncover patterns, identify emerging risks, and develop predictive models for assessing claim probabilities, estimating aggregate claim amounts, and deriving necessary optimal level of reserves to set aside to cover these future estimated claims. See Klugman et al., (2019)
@@
The two-part frequency-severity models have historically been the norm used in insurance loss data analytics for modeling aggregate claims for pure premium. These models recognize that the frequency of claims (how often claims occur) and the severity of claims (the amount paid per claim) are often influenced by different factors and follow different distributions. The claim frequency model focuses on predicting the number of claims that occur within a specified period, such as a year or a month, and often follows a count distribution model, such as Poisson or negative binomial distribution. It considers factors that influence claim occurrence, such as policyholder characteristics (e.g., age, location), policy features (e.g., coverage type, deductible), external factors (e.g., weather conditions, economic trends), and any relevant historical data. On the other hand, the severity model focuses on predicting the monetary value of individual claims, i.e., how much each claim will cost the insurer. It generally follows a continuous distribution, such as gamma or log-normal, that captures the variability in claim amounts and their right-skewed nature. It also considers factors that may influence claim amounts. See Frees et al., 2014a .
@@
Since the introduction of Tweedie et al., (1984) in the 1980s, the Tweedie mixture distribution has gained popularity in insurance claims modeling as it simplifies the modeling process by eliminating the need for separate frequency and severity models. The distribution itself exhibits a point mass at zero representing the probability of zero claims, followed by continuous density function for positive claim values. See Frees et al., 2014a . Such integration is a more accurate risk assessment as it also captures the inherent interdependence between the frequency and severity components. Primarily due to the distribution’s exceptional flexibility to accommodate a diverse range of data, Tweedie regression models have a proven track record of providing superior fits to historical insurance claims data when compared to the conventional two-part models.
@@
With the influx of machine learning, the abovementioned conventional methods started to adopt powerful boosting algorithms. Gradient boosting has become immensely popular in the field of machine learning and is increasingly being adopted in insurance and actuarial science for many reasons, among which include predictive accuracy and robustness. This technique was initially introduced by Friedman, (2001) and is based around the idea of combining multiple weak learners, such as small decision trees, to create a single, strong, and robust predictive model. Suppose we are given a training dataset $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}^{n}_{1}$. Gradient boosting generates a sequence of functions  $W_{0},W_{1},\cdots,W_{T}$, by minimizing the expected value of a specified loss function, $\ell(y_{i},W_{t})$. This loss function takes two inputs: the $i$-th response $y_{i}$ and the $t$-th function $W_{t}$, which estimates $y_{i}$. To improve on these estimates of $y_{i}$, we iteratively search for another function $W_{t+1}=W_{t}+w_{t+1}(\bm{x})$ that minimizes the expected value of the loss function, where $w_{t+1}$ is the weak learner at the $(t+1)$-th iteration. That is,
@@
Our goal is to minimize the value of the loss function with respect to $y$ and $W_{t}$, while considering the additional factor $w_{t+1}$. Given that the loss function is continuous and differentiable, this process involves adjusting $W_{t}$ towards the direction where the loss function decreases most steeply. This direction corresponds to the rate of change of the loss function. Thus, by aligning $w_{t+1}$ with the direction where the gradient of the loss function with respect to $W_{t}$ declines most sharply, we approximate a $w_{t+1}$ that closely approaches the minimum of $\sum_{i=1}^{n}\ell(y_{i},W_{t+1}(\bm{x}_{i}))$. The direction of steepest descent in this context is the negative gradient. Consequently, we define the pseudo-residuals $r_{i,t+1}$ as follows:
@@
Since we are interested in modeling aggregate insurance claims, we design boosting algorithms to optimize loss functions on the distribution of the target variable, e.g., Tweedie loss function calculated as the negative log-likelihood of the data. Insurance claims data frequently features a significant number of zero values, especially for low-frequency events or claims. Additionally, the data often includes a small number of high-value claims, leading to a skewed distribution with extreme values. Given these inherent characteristics of insurance claims data, choosing a loss function associated with the Tweedie distribution is a promising strategy when incorporating gradient boosting methods. In the case of the Tweedie loss function, it incorporates three parameters $\mu$, $p$, and $\phi$ that all control the shape of the function. We find from Figure 1(a) the lack of flexibility of these parameters to control the function in handling the zeros and right-skewness of the data.
Henceforth, we propose an enhancement to the Tweedie distribution through the introduction of an inflation probability, denoted as $q$. This modification leads to the formulation of what we refer to as the Zero-Inflated Tweedie distribution. In this revised distribution, the probability of zero is represented by the expression $(q+(1-q)\times f_{Tw}(y=0)$, where $f_{Tw}(y=0)$ denotes the probability of a zero from the original Tweedie distribution. In order to provide an even further flexibility in accommodating the inherent characteristics of the data, we suggest introducing the inflation probability $q$ as a function dependent on $\mu$, with an additional parameterization $\gamma$. Figure 1(b) showcases a more flexible $q$ as a function of $\mu$, addressing the limitations of the Tweedie distribution.
@@
To enhance the flexibility of modeling insurance losses, we made adjustments to the conventional gradient boosted trees approach in the following manners. First, we implemented a customized loss function based on the zero-inflated Tweedie distribution described above. This decision was made because we believe the traditional Tweedie distribution is insufficient to provide a more accurate characterization of insurance losses. While there are some papers (cite come) that discuss its potential, the use of zero-inflated Tweedie distribution remains limited in the actuarial and statistical literature. Second, we next reparameterized the zero-inflated Tweedie loss function to enable the modeling of the inflation probability $q$ as a function of the mean parameter $\mu$. Because of this implied relationship between $q$ and $m$, this results in a singular model that becomes simplified for purposes of model interpretation and model efficiency. Such approach becomes more attractive to practitioners and hence, becomes highly likely to be adopted in practice. In the data analysis section, we demonstrated various ways to interpret the model’s estimates and outcomes. Third, for empirical demonstration, we analyzed a synthetic telematics dataset drawn from So et al., (2021). This data contains records of 100,000 automobile insurance policies and showcases a notable zero-inflation pattern, with only 2,698 policies experiencing at least one claim. We specifically picked this dataset to allow us to examine the need for any specific allowance or special treatments related to compositional features. It is known that when GLM (Generalized Linear Model) is used, we need adjustments to the compositional feature variables to avoid bias. However, we found that such adjustments become unnecessary when boosting algorithms are employed. Finally, we implemented CatBoost, among other boosting libraries, to train our Gradient Boosted Decision Tree (GBDT) models. In So, (2024), So explored various boosting libraries, including XGBoost, LightGBM, and CatBoost, on the same dataset for estimating various frequency models and found that CatBoost is the most efficient. For our purposes, CatBoost further offers enhanced efficiency in handling special types of feature variables, particularly categorical features that are not uncommon in insurance data.
By leveraging the zero-inflated Tweedie loss function and employing CatBoost, we were able to create a more flexible and accurate model for predicting insurance losses. These adjustments not only improved the model’s performance but also enhanced its interpretability and ability to handle real-world insurance data effectively.
@@ 2 Zero-Inflated Tweedie Boosted Decision Trees
$Y=\begin{cases}0,&\text{if }N=0,\\[4.30554pt] X_{1}+X_{2}+\cdots+X_{N},&\text{if }N\geq 1.\end{cases}$ (4)
@@ 3 Categorical Boosting: CatBoost
Boosting algorithms have gained immense popularity among actuaries as they are considered one of the most advanced learning algorithms developed in recent years. This is because boosting is an iterative fitting process that combines weak learners – those slightly better than random – into a strong learner, resulting in more precise and accurate predictions. This process is often referred to as a stage-wise additive model, where a new weak learner is added at a time while existing weak learners in the model remain fixed and unchanged. (need actuarial boosting papers: Michel Denuit, check out Montse Guillen)
@@
CatBoost, developed by Yandex ((Prokhorenkova et al.,, 2018)), is recognized for its effectiveness in handling heterogeneous datasets, a common scenario in insurance data. Recent studies (So, (2024)) have demonstrated CatBoost’s superior performance compared to its counterparts when processing insurance data. This section will elaborate on CatBoost’s distinctive implementation strategy.
@@
Catboost stands out among GBDT libraries due to its unique handling of categorical variable. It is important to note that categorical features do not inherently possess a natural ordering, and therefore, before making a split in the decision tree, categorical features must be converted into numerical values. CatBoost addresses this challenge by employing a technique known as “Ordered Target Statistic.” This method involves randomly rearranging the dataset and then performing target encoding for each instance, taking into account only the instances that precede the current one. The process of transforming categorical features into their numerical counterparts in CatBoost involves the following steps:
@@
It is essential to ensure that the target encoding is solely based on instances preceding the current one to prevent data leakage. Consequently, both “countInClass” and “totalCount” are computed following this rule.
@@ 4.1 Scenario 1: Functionally unrelated
$\ell(W^{mean}_{T}(\bm{x}),W^{prob}_{T}(\bm{x}))=\begin{cases}-\ln\left(q+(1-q)\cdot\exp\left(-\displaystyle\frac{1}{\phi}\displaystyle\frac{\mu^{2-p}}{2-p}\right)\right),&\quad\text{if }y=0\\[12.91663pt] -\ln(1-q)-\ln a(y,\phi,p)-\left(\displaystyle\frac{1}{\phi}\left(y\cdot\frac{\mu^{1-p}}{1-p}-\frac{\mu^{2-p}}{2-p}\right)\right),&\quad\text{if }y>0,\\ \end{cases}$ (18)
where $\mu=Ee^{W^{mean}_{T}}$ and $q=\text{logit}^{-1}(W^{prob}_{T}(\bm{x}))$. The gradients and Hessians of the loss function with respect to $W^{mean}_{t}$ are given by:
$g^{mean}_{t+1}=\partial_{W^{mean}_{t}}\ell(y,W^{mean}_{t},W^{prob}_{t})=\begin{cases}\displaystyle\frac{\alpha\cdot\beta}{q+\alpha},&\quad\text{if }y=0\\[12.91663pt] -\displaystyle\frac{1}{\phi}y\mu^{1-p}+\beta,&\quad\text{if }y>0,\\ \end{cases}$ (19)
$h^{mean}_{t+1}=\partial^{2}_{W^{mean}_{t}}\ell(y,W^{mean}_{t},W^{prob}_{t})=\begin{cases}\displaystyle\frac{\alpha\cdot\beta((2-p-\beta)\cdot(q+\alpha)+\alpha\cdot\beta)}{(q+\alpha)^{2}},&\quad\text{if }y=0\\[12.91663pt] -\displaystyle\frac{1}{\phi}y(1-p)\mu^{1-p}+(2-p)\cdot\beta,&\quad\text{if }y>0,\\ \end{cases}$ (20)
@@
Similarly, the gradients and the Hessians of the loss function with respect to $W^{prob}_{t}$ are given by:
$g^{prob}_{t+1}=\partial_{W^{prob}_{t}}\ell(y,W^{mean}_{t+1},W^{prob}_{t})=\begin{cases}\displaystyle\frac{1}{1+e^{\delta+W^{logit}_{t}}}-\frac{1}{1+e^{W^{logit}_{t}}},&\quad\text{if }y=0\\[12.91663pt] \displaystyle\frac{e^{W^{logit}_{t}}}{1+e^{W^{logit}_{t}}},&\quad\text{if }y>0,\\ \end{cases}$ (21)
$h^{prob}_{t+1}=\partial^{2}_{W^{prob}_{t}}\ell(y,W^{mean}_{t+1},W^{prob}_{t})=\begin{cases}\displaystyle\frac{e^{W^{prob}_{t}}}{\left(1+e^{W^{prob}_{t}}\right)^{2}}-\frac{e^{\delta+W^{prob}_{t}}}{\left(1+e^{\delta+W^{prob}_{t}}\right)^{2}},&\quad\text{if }y=0\\[12.91663pt] \displaystyle\frac{e^{W^{prob}_{t}}}{\left(1+e^{W^{prob}_{t}}\right)^{2}},&\quad\text{if }y>0,\\ \end{cases}$ (22)
where $\mu=Ee^{W^{mean}_{t+1}}$ and $\delta=\displaystyle\frac{1}{\phi}\displaystyle\frac{\mu^{2-p}}{2-p}$.
To minimize the loss function with two separate models, we adopt the training technique proposed by Meng et al., (2022). This approach employs the principle of coordinate descent optimization, where during each iteration, one parameter (e.g., $\mu$ or $q$) is optimized while keeping the other fixed. This process is repeated alternately for $\mu$ and $q$ until convergence, leading to improve predictive accuracy for both parameters. More specifically, during the $t$-th iteration, with a fixed inflation parameter $q$, we train $w_{t}^{\text{mean}}$ (the $t$-th decision tree) to estimate the mean parameter $\mu$. Then, with the estimated $\mu$ held constant, we train $w_{t}^{\text{prob}}$ to estimate the inflation probability $q$. After completion of an estimation cycle for both $\mu$ and $q$, the dispersion parameter $\phi$ is estimated by minimizing the mean deviation described in subsection 5.1.1. The power parameter is set to a predetermined constant, typically 1.5. For a detailed explanation of this training algorithm, please refer to Algorithm 1 in Appendix A.

### v2

@@ Abstract
In this paper, we explore advanced modifications to the Tweedie regression model in order to address its limitations in modeling aggregate claims for various types of insurance such as automobile, health, and liability. Traditional Tweedie models, while effective in capturing the probability and magnitude of claims, usually fall short in accurately representing the large incidence of zero claims. Our recommended approach involves a refined modeling of the zero-claim process, together with the integration of boosting methods in order to help leverage an iterative process to enhance predictive accuracy. Despite the inherent slowdown in learning algorithms due to this iteration, several efficient implementation techniques that also help precise tuning of parameters like XGBoost, LightGBM, and CatBoost have emerged. Nonetheless, we chose to utilize CatBoost, an efficient boosting approach that effectively handles categorical and other special types of data. The core contribution of our paper is the assembly of separate modeling for zero claims and the application of tree-based boosting ensemble methods within a CatBoost framework, assuming that the inflated probability of zero is a function of the mean parameter. The efficacy of our enhanced Tweedie model is demonstrated through the application of an insurance telematics dataset, which presents the additional complexity of compositional feature variables. Our modeling results reveal a marked improvement in model performance, showcasing its potential to deliver more accurate predictions suitable for insurance claim analytics.
Keywords: Tweedie, zero-inflated models, tree-based boosting, gradient boosting, CatBoost, compositional features, aggregate loss models.
@@ 1 Introduction
Insurance loss analytics, which is the backbone of modern insurance risk management strategies, involves a systematic collection, processing, and analysis of observed insurance loss data to extract insights and produce predictive models to help quantify and manage risk exposure. Actuaries and risk managers have historically employed the use of data analytics to uncover patterns, identify emerging risks, and develop predictive models for assessing claim probabilities, estimating aggregate claim amounts, and deriving necessary optimal level of reserves to set aside to cover these future estimated claims. See Klugman et al., (2019).
@@
The two-part frequency-severity models have historically been the norm used in insurance loss data analytics for modeling aggregate claims for pure premium. These models recognize that the frequency of claims (how often claims occur) and the severity of claims (the amount paid per claim) are often influenced by different factors and follow different distributions. The claim frequency focuses on predicting the number of claims that occur within a specified period, such as a year or a month, and often follows a count distribution model, such as Poisson or negative binomial distribution. It considers factors that influence claim occurrence, such as policyholder characteristics (e.g., age, location), policy features (e.g., coverage type, deductible), external factors (e.g., weather conditions, economic trends), and any relevant historical data. On the other hand, the severity focuses on predicting the monetary value of individual claims, i.e., how much each claim will cost the insurer. It generally follows a continuous distribution, such as gamma or log-normal, that captures the variability in claim amounts and their right-skewed nature. It also considers factors that may influence claim amounts. See Frees et al., 2014a .
@@
Since the introduction of Tweedie et al., (1984) in the 1980s, the Tweedie mixture distribution has gained popularity in insurance claims modeling as it simplifies the modeling process by eliminating the need for separate frequency and severity models. The distribution itself exhibits a point mass at zero representing the probability of zero claims, followed by continuous density function for positive claim values. See Frees et al., 2014a . Such integration is a more accurate risk assessment as it also captures the inherent interdependence between the frequency and severity components. Primarily due to the distribution’s exceptional flexibility to accommodate a diverse range of data, Tweedie regression models have a proven track record of providing superior fits to historical insurance claims data when compared to the conventional two-part models. It is true that the two-part frequency-severity models can accommodate ease of implementation of excess of zeros (in the frequency part) and can leverage modern machine learning to enhance prediction (in the severity component), e.g., Shi et al., (2024). However, this poses a significant drawback of being able to build a unified framework to allow us to deduce practical interpretation and understanding of the effect of the features affecting the aggregate loss, as we illustrate in the latter part of this paper.
@@
With the influx of machine learning, the abovementioned conventional methods started to adopt powerful boosting algorithms. Gradient boosting has become immensely popular in the field of machine learning and is increasingly being adopted in insurance and actuarial science for many reasons, among which include predictive accuracy and robustness. This technique was initially introduced by Friedman, (2001) and is based around the idea of combining multiple weak learners, such as small decision trees, to create a single, strong, and robust predictive model. Suppose we are given a training dataset $\mathcal{D}=\{(\bm{x}_{i},y_{i})\}^{n}_{1}$. Gradient boosting generates a sequence of functions  $W_{0},W_{1},\ldots,W_{T}$, by minimizing the expected value of a specified loss function, $\ell(y_{i},W_{t})$. This loss function takes two inputs: the $i$-th response $y_{i}$ and the $t$-th function $W_{t}$, which estimates $y_{i}$. To improve on these estimates of $y_{i}$, we iteratively search for another function $W_{t+1}=W_{t}+w_{t+1}(\bm{x})$ that minimizes the expected value of the loss function, where $w_{t+1}$ is the weak learner at the $(t+1)$-th iteration. That is,
@@
Our goal is to minimize the value of the loss function with respect to $y$ and $W_{t}$, while considering the additional factor $w_{t+1}$. Assuming that the loss function is continuous and differentiable, this process involves adjusting $W_{t}$ towards the direction where the loss function decreases most steeply. This direction corresponds to the rate of change of the loss function. Thus, by aligning $w_{t+1}$ with the direction where the gradient of the loss function with respect to $W_{t}$ declines most sharply, we approximate a $w_{t+1}$ that closely approaches the minimum of $\sum_{i=1}^{n}\ell(y_{i},W_{t+1}(\bm{x}_{i}))$. The direction of steepest descent in this context is the negative gradient. Consequently, we define the pseudo-residuals $r_{i,t+1}$ as follows:
@@
Since we are interested in modeling aggregate insurance losses, we design boosting algorithms to optimize loss functions on the distribution of the target variable, e.g., Tweedie loss function calculated as the negative log-likelihood of the data. Insurance claims data frequently features a significant number of zero values, especially for low-frequency events or claims. Additionally, the data often includes a small number of high-value claims, leading to a skewed distribution. Given these inherent characteristics of insurance claims data, choosing a loss function associated with the Tweedie distribution is a promising strategy when incorporating gradient boosting methods. In the case of the Tweedie loss function, it incorporates three parameters $\mu$, $p$, and $\phi$ that all control the shape of this function. We find from Figure 1(a) the lack of flexibility of these parameters to control the function in handling the zeros and right-skewness of the data.
Henceforth, we propose an enhancement to the Tweedie distribution through the introduction of an inflation probability, denoted as $q$. This modification leads to the formulation of what we refer to as the Zero-Inflated Tweedie distribution. In this revised distribution, the probability of zero is represented by the expression $(q+(1-q)\times f_{Tw}(y=0))$, where $f_{Tw}(y=0)$ denotes the probability of a zero from the original Tweedie distribution. In order to provide an even further flexibility in accommodating the inherent characteristics of the data, we suggest introducing the inflation probability $q$ as a function dependent on $\mu$, with an additional parameterization $\gamma$. Figure 1(b) showcases a more flexible $q$ as a function of $\mu$, addressing the limitations of the Tweedie distribution.
@@
To address the challenges of zero-inflated, skewed data and to enhance the flexibility of modeling insurance losses, our paper makes several key contributions. First, we implemented a customized loss function based on the zero-inflated Tweedie distribution described above. This decision was made because we believe the traditional Tweedie distribution is insufficient to provide a more accurate characterization of insurance losses. While there are some papers that discuss its potential, the use of zero-inflated Tweedie distribution remains limited in the actuarial and statistical literature. Second, we reparameterized the zero-inflated Tweedie loss function to model the inflation probability $q$ as a function of the mean parameter $\mu$. This relationship between $q$ and $\mu$ results in a unified model that simplifies interpretation and improves computational efficiency. This approach is more appealing to practitioners, increasing the likelihood of its adoption in practice. In the data analysis section, we demonstrated various methods for interpreting the model’s estimates and outcomes, which are made possible by our unified model. In addition, in contrast to the standard Tweedie model where frequency and severity are assumed independent, we allow for a direct link between $q$ and $\mu$. This implies some form of dependence or correlation between loss frequency and severity, which makes the model much more flexible as indicated by Shi and Zhao, (2020). Third, for empirical demonstration, we analyzed a synthetic telematics dataset drawn from So et al., (2021). This data contains records of 100,000 automobile insurance policies and showcases a notable zero-inflation pattern, with only 2,698 policies experiencing at least one claim. We specifically picked this dataset to allow us to examine the need for any specific allowance or special treatments related to compositional features. It is known that when GLM (Generalized Linear Model) is used, we need adjustments to the compositional feature variables to avoid bias. However, we found that such adjustments become unnecessary when boosting algorithms are employed. Finally, we implemented CatBoost, among other boosting libraries, to train our Gradient Boosted Decision Tree (GBDT) models. In So, (2024), So explored various boosting libraries, including XGBoost, LightGBM, and CatBoost, on the same dataset for estimating various frequency models and found that CatBoost is the most efficient. For our purposes, CatBoost further offers enhanced efficiency in handling special types of feature variables, particularly categorical features that are not uncommon in insurance data. By leveraging the zero-inflated Tweedie loss function and employing CatBoost, we were able to create a more flexible and accurate model for predicting insurance losses. These adjustments not only improved the model’s performance but also enhanced its interpretability and ability to handle real-world insurance data effectively.
@@ 2 Zero-Inflated Tweedie Boosted Decision Trees
$Y=\begin{cases}0,&\text{if }N=0,\\[4.30554pt] X_{1}+X_{2}+\cdots+X_{N},&\text{if }N=1,2,\ldots.\end{cases}$ (4)
@@ 3 Categorical Boosting: CatBoost
Boosting algorithms have gained immense popularity among actuaries as they are considered one of the most advanced learning algorithms developed in recent years. This is because boosting is an iterative fitting process that combines weak learners – those slightly better than random – into a strong learner, resulting in more precise and accurate predictions. This process is often referred to as a stage-wise additive model, where a new weak learner is added at a time while existing weak learners in the model remain fixed and unchanged.
@@
CatBoost, developed by Yandex (Prokhorenkova et al.,, 2018), is recognized for its effectiveness in handling heterogeneous datasets, a common scenario in insurance data. Recent studies (So, (2024)) have demonstrated CatBoost’s superior performance compared to its counterparts when processing insurance data. This section will elaborate on CatBoost’s distinctive implementation strategy.
@@
Catboost stands out among GBDT libraries due to its unique handling of categorical variable. Other libraries such as XGBoost and LightGBM do not directly support, but need some preprocessing, in handling categorical data. It is important to note that categorical features do not inherently possess a natural ordering, and therefore, before making a split in the decision tree, categorical features must be converted into numerical values. CatBoost addresses this challenge by employing a technique known as “Ordered Target Statistic.” This method involves randomly rearranging the dataset and then performing target encoding for each instance, taking into account only the instances that precede the current one. The process of transforming categorical features into their numerical counterparts in CatBoost involves the following steps:
@@
It is essential to ensure that the target encoding is solely based on instances preceding the current one to prevent data leakage. Consequently, both “countInClass” and “totalCount” are computed following this rule. For further details, please see Hancock and Khoshgoftaar, (2020).
@@ 4.1 Scenario 1: Functionally unrelated
$\ell(y,W^{mean}_{T}(\bm{x}),W^{prob}_{T}(\bm{x}))=\begin{cases}-\ln\left(q+(1-q)\cdot\exp\left(-\displaystyle\frac{1}{\phi}\displaystyle\frac{\mu^{2-p}}{2-p}\right)\right),&\quad\text{if }y=0\\[12.91663pt] -\ln(1-q)-\ln a(y,\phi,p)-\left(\displaystyle\frac{1}{\phi}\left(y\cdot\frac{\mu^{1-p}}{1-p}-\frac{\mu^{2-p}}{2-p}\right)\right),&\quad\text{if }y>0,\\ \end{cases}$ (18)
where $\mu=Ee^{W^{mean}_{T}(\bm{x})}$ and $q=\text{logit}^{-1}(W^{prob}_{T}(\bm{x}))$. The gradients and Hessians of the loss function with respect to $W^{mean}_{t}(\bm{x})$ are given by:
$g^{mean}_{t+1}=\partial_{W^{mean}_{t}(\bm{x})}\ell(y,W^{mean}_{t}(\bm{x}),W^{prob}_{t}(\bm{x}))=\begin{cases}\displaystyle\frac{\alpha\cdot\beta}{q+\alpha},&\quad\text{if }y=0\\[12.91663pt] -\displaystyle\frac{1}{\phi}y\mu^{1-p}+\beta,&\quad\text{if }y>0,\\ \end{cases}$ (19)
[…]
@@
Similarly, the gradients and the Hessians of the loss function with respect to $W^{prob}_{t}(\bm{x})$ are given by:
$g^{prob}_{t+1}=\partial_{W^{prob}_{t}(\bm{x})}\ell(y,W^{mean}_{t+1}(\bm{x}),W^{prob}_{t}(\bm{x}))=\begin{cases}\displaystyle\frac{1}{1+e^{\delta+W^{prob}_{t}(\bm{x})}}-\frac{1}{1+e^{W^{prob}_{t}(\bm{x})}},&\quad\text{if }y=0\\[12.91663pt] \displaystyle\frac{e^{W^{prob}_{t}(\bm{x})}}{1+e^{W^{prob}_{t}(\bm{x})}},&\quad\text{if }y>0,\\ \end{cases}$ (21)
$h^{prob}_{t+1}=\partial^{2}_{W^{prob}_{t}(\bm{x})}\ell(y,W^{mean}_{t+1}(\bm{x}),W^{prob}_{t}(\bm{x}))=\begin{cases}\displaystyle\frac{e^{W^{prob}_{t}(\bm{x})}}{\left(1+e^{W^{prob}_{t}(\bm{x})}\right)^{2}}-\frac{e^{\delta+W^{prob}_{t}(\bm{x})}}{\left(1+e^{\delta+W^{prob}_{t}(\bm{x})}\right)^{2}},&\quad\text{if }y=0\\[12.91663pt] \displaystyle\frac{e^{W^{prob}_{t}(\bm{x})}}{\left(1+e^{W^{prob}_{t}(\bm{x})}\right)^{2}},&\quad\text{if }y>0,\\ \end{cases}$ (22)
where $\mu=Ee^{W^{mean}_{t+1}(\bm{x})}$ and $\delta=\displaystyle\frac{1}{\phi}\displaystyle\frac{\mu^{2-p}}{2-p}$.
To minimize the loss function with two separate models, we adopt the training technique proposed by Meng et al., (2022). This approach employs the principle of coordinate descent optimization, where during each iteration, one parameter (e.g., $\mu$ or $q$) is optimized while keeping the other fixed. This process is repeated alternately for $\mu$ and $q$ until convergence, leading to improve predictive accuracy for both parameters. More specifically, during the $t$-th iteration, with a fixed inflation parameter $q$, we train $w_{t}^{mean}(\bm{x})$ (the $t$-th decision tree) to estimate the mean parameter $\mu$. Then, with the estimated $\mu$ held constant, we train $w_{t}^{prob}(\bm{x})$ to estimate the inflation probability $q$. After completion of an estimation cycle for both $\mu$ and $q$, the dispersion parameter $\phi$ is estimated by minimizing the mean deviation described in subsection 5.1.1. The power parameter is set to a predetermined constant, typically 1.5. For a detailed explanation of this training algorithm, please refer to Algorithm 1 in Appendix A.